Consider a homogeneous product industry with inverse demand function p-35 -Q a) Assume that the industry is initially monopolized by an incumbent firm (firm I) which has the exclusive right to use the state-of-the-art technology summarized by the total cost function C-10q. Find the initial monopoly equilibrium (price, quantity, industry profit, consumer surplus and total surplus) and the associated degrees of concentration (Herfindahl index) and market power (Lerner index) b) Assume now that a new firm (firm N) discovers and patents a more efficient technology, summarized by the total cost function C-8q. The new technology can be used only by the new firm, which enters the market and compete in prices with the previous monopolist (according to the Bertrand model with asymmetric costs). Find the resulting Bertrand equilibrium (price, quantity, industry profit, consumer surplus and total surplus) and the associated degrees of concentration (Herfindahl index) and market power (Lerner index)
Answer #1

p= 35-q is an inverse demand function .............(1)

a) A monopoly firm is in equilibrium whenever its marginal cost(MC) equals to Marginal revenue(MR)

We have given Total Cost(C) = 10q

=> Differentiating it with respect to 'q' we can find marginal cost of Monopoly firm

i.e.   

10orMC10

For marginal Revenue, Differentiating total revenue(TR) with respect to 'q'

TR = p*q

=> (35-q)*q

TR= 35q-q2

Differentiating this euation with respect to 'q' we get

TR).2 = 35-2qorMR = 35-2q

Equating MR=MC we get the equilibrium Quantity i.e.

MR=MC

35-2q=10

=>q=12.5 ...........................(2)

=>At q=12.5 By equation (1)

'p' equals to 22.5 ...........................(3)

Difference between Total Revenue and Total Cost is industry's profit(prod)

II TR-TO (TR= 35*12.5 - 12.52 and TC=10*12.5)

281.25 125 -

156.25 ......................(4)

Consumer Surplus* (35- P)

=> 1/2* 22.5*22.5

CS= 253.125 .........................(5)

ProducerSur plus = (P-M R) * q +-* (M R-MC) * q

PS = (22.5-10)*12.5 + 1/2*(10-10)*12.5

PS= 156.25 .........................(6)

Total Surplus = Consumer Surplus + Producr Surplus

=> 253.125+156.25

TS= 409.375 ..................(7)

There is only single firm in this monopoly situation which means market share of this firm is 100%, therefore Herfindahl Index = 1002=10,000.

Lerner Index equals

L=rac{p-MC}{p}

22.5-10 う? 5

Lerner Index = 55.6% ................(8)

b) In Bertrand Competition, Comparing MCs, we can see that MC1=10 and MC2=8  

So MC1>MC2

then Bertrand equilibrium implies that p* = MC1 which means firm 2 will capture all the market

Therefore firm 1's profit equal to 0

Firm 2's profit equals (MC1-MC2)*Q(MC1)

=> (10-8)*25

50

Total Profit = 50

Herfindahl Index for firm 1 will be 0 and for firm 2 will b 10,000

Lerner Index for firm 2 will be (10 - 8)/10

=> 20%